The First Law of Thermodynamics

A gaseous reaction occurs at a constant pressure of 50.0 atm and releases 64.0 kJ of heat. Before the reaction, the volume of the system was 6.60 L. After thereaction, the volume of the system was 3.00 L. Calculate the total internal energy change, Delta U, in kilojoules.
Answer #1

Δq =ΔU+PΔV

P= 50 atm

ΔQ= 64KJ/mol

ΔU = Δq-pΔV = -64.0-50(3-6.60) = 101 KJ/mol = change in internal energy

answered by: Erensto